Assume the​ following: Melita carried an average daily balance of $937 on her credit card this month. Her previous balance last month was ​$1,717​, compared to a balance of ​$227 this month. There are 30 days in this billing cycle and Melita always makes a payment on the fifteenth of the month. Based on this​ information, calculate the monthly interest charges for credit card accounts charging 14 ​percent, 16 ​percent, and 18 percent interest. Complete the following chart. Since the average daily balance is the most commonly used balance calculation​ method, is shopping for a lower interest rate really that​ important?

14​%

16​%

18​%

Average Daily Balance

​$10.93

Previous Balance

​$22.89

Adjusted Balance

​$3.41

The three primary methods used to determine interest charges on unpaid credit balances are​ (1) the average daily balance​ method, (2) the previous balance​ method, and​ (3) the adjusted balance method.
Answer #1

Average daily balance interest = Average daily balance*Rate/12

14% interest= 937*14%/12 = 10.93

16% interest = 937*16%/12=12.49

18% interest = 937*18%/12 = 14.06

Previous balance interest = Previous balance*Rate/12

14% interest= 1717*14%/12 = 20.03

16% interest = 1717*16%/12=22.89

18% interest = 1717*18%/12 = 25.76

Adjusted balance interest= Adjusted balance* Rate/12

14% interest= 227*14%/12 = 2.65

16% interest = 227*16%/12= 3.03

18% interest = 227*18%/12 = 3.41